OpenStudy (anonymous):

Does anyone here go to k12?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

This is where I'm getting my questions from.

OpenStudy (acxbox22):

please plose this question it will be better if u ask in the chats posts like these are considered spam thank you ;)

OpenStudy (alexandervonhumboldt2):

no matter please post question

OpenStudy (acxbox22):

close*

OpenStudy (alexandervonhumboldt2):

@acxbox22 tht probably a question some users want help form same school

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Do any of you know how to do graph homework? or how to find equations with graphs???

OpenStudy (anonymous):

Hello? @AlexandervonHumboldt2 ? @acxbox22 ?

OpenStudy (acxbox22):

yes post your question

OpenStudy (anonymous):

:S fine

OpenStudy (anonymous):

Choose the correct answer. Elena bought 320 tropical fish for a museum display. She bought 7 times as many triggerfish as parrotfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of parrotfish Elena bought and y represents the number of triggerfish she bought? A. x + y = 7 y = 320x B. x + 7y = 320 y = 7x C. x + y = 320 y = 7x D. x – y = 320 y = 7x Can you help me work this out??

OpenStudy (anonymous):

..... (2 years later)

OpenStudy (acxbox22):

she bought a total of 320 fish right let x be the amount of triggerfish and y parrotfish from this we know the sum of x and y equals 320

OpenStudy (anonymous):

It say 7 time though... so is the trigger and parrotfish 7?

OpenStudy (acxbox22):

x+y=320 is our first equation since the number of both fish combine is 320 our second equation will be 7 times the number of triggerfish(x) which will equal the number of parrotfish

OpenStudy (anonymous):

Oh.... well lemme try doing it

OpenStudy (anonymous):

okay is it x + y = 320y = 7x?